# A Unified Implementation of Quasi-Monte Carlo Generators, Randomization Routines, and Fast Kernel Methods

### v1

@@ 4.2. Digital nets
and ${\boldsymbol{\Delta}}_{j}$ is the $j^{\mathrm{th}}$ column of ${\mathsf{\Delta}}$. Randomly shifted digital nets use ${\mathsf{\Delta}}\overset{\mathrm{IID}}{\sim}{\mathcal{U}}\{0,\dots,b-1\}^{t_{\mathrm{max}}\times d}$ i.e. is each digit is chosen uniformly from $\{0,\dots,b-1\}$.
Digitally permuted digital net:
In what follows will denote permutations of $\{0,\dots,b-1\}$ by $\pi$. Suppose we are given a set of permutations
@@
Then we may construct the digitally permuted digital net $\{{\boldsymbol{x}}_{i}\}_{i=1}^{b^{m}-1}$ where
@@
Nested Uniform Scrambling (NUS):
NUS is often called Owen scrambling for its conception in (Owen, 1995). As before, $\pi$ values denote permutations of $\{0,\dots,b-1\}$. Now suppose we are given a set of permutations
@@
Then a nested uniform scrambling of a digital net is $\{{\boldsymbol{x}}_{i}\}_{i=1}^{b^{m}-1}$ where
@@
our implementation cannot afford to generated all permutations apriori. Instead, permutations are generated and stored only as needed. As with digitally-permuted digital nets, NUS uses independent uniform random permutations.
The randomization routines described above are visualized in fig. 2.
@@
for $j=1,\dots,d$. Following (Owen, 2003), let us denote elements in $\{1,\dots,b\}$ by $h$ and elements in $\{0,\dots,b\}$ by $g$. Then common structures for ${\mathsf{S}}_{j}$ include
@@
which corresponds to Matousek’s linear scrambling (Matoušek, 1998), Tezuka’s $i$-binomial scrambling (Tezuka, 2002), and Owen’s striped LMS (Owen, 2003) (not to be confused with NUS which is often called Owen scrambling). Random LMS chooses $g$ and $h$ values all independently and uniformly from $\{1,\dots,b\}$ and $\{0,\dots,b-1\}$ respectively. These scrambling are described in more detail and connected to NUS in (Owen, 2003).
Digital interlacing enables the construction of higher-order digital nets. Denote the digitally interlaced version of ${\mathsf{C}}_{1},{\mathsf{C}}_{2},\dots\in\{0,\dots,b\}^{t_{\mathrm{max}}\times m}$ with interlacing factor $\alpha\in{\mathbb{N}}$ by ${\widetilde{\mathsf{C}}}_{1},{\widetilde{\mathsf{C}}}_{2},\dots\in\{0,\dots,b\}^{\alpha t_{\mathrm{max}}\times m}$. Then we have ${\widetilde{\mathsf{C}}}_{jtk}=C_{{\widetilde{j}},\tt,k}$ where ${\widetilde{j}}=\alpha(j-1)+(t\mod\alpha)+1$ and $\tt=\lfloor t/\alpha\rfloor$ for $j\geq 1$ and $0\leq t<\alpha t_{\mathrm{max}}$ and $1\leq k\leq m$. For example, with $m=2$, $t_{\mathrm{max}}=2$, and $\alpha=2$ we have the following
${\mathsf{C}}_{1}=\begin{pmatrix}c_{101}&c_{102}\\ c_{111}&c_{112}\end{pmatrix}\qquad{\mathsf{C}}_{2}=\begin{pmatrix}c_{201}&c_{202}\\ c_{211}&c_{212}\end{pmatrix}\qquad{\mathsf{C}}_{3}=\begin{pmatrix}c_{301}&c_{302}\\ c_{311}&c_{312}\end{pmatrix}\qquad{\mathsf{C}}_{4}=\begin{pmatrix}c_{401}&c_{402}\\ c_{411}&c_{412}\end{pmatrix}\qquad\dots$
${\widetilde{\mathsf{C}}}_{1}=\begin{pmatrix}c_{101}&c_{102}\\ c_{201}&c_{202}\\ c_{111}&c_{112}\\ c_{211}&c_{212}\end{pmatrix}\qquad{\widetilde{\mathsf{C}}}_{2}=\begin{pmatrix}c_{301}&c_{302}\\ c_{401}&c_{402}\\ c_{311}&c_{312}\\ c_{411}&c_{412}\end{pmatrix}\qquad\dots$
Higher-order NUS requires NUS be applied to the digital net point sets generated by ${\mathsf{C}}_{1},\dots,{\mathsf{C}}_{\alpha d}$ and then interlacing be performed to the digits $\{{\boldsymbol{\mathsf{z}}}_{i1}\}_{i=0}^{b^{m}-1},\dots,\{{\boldsymbol{\mathsf{z}}}_{i,\alpha d}\}_{i=0}^{b^{m}-1}$ (Dick, 2011). For higher-order LMS, we apply LMS to the generating matrices ${\mathsf{C}}_{1},\dots,{\mathsf{C}}_{\alpha d}$, then interlace them, then generate the digital net. As we shown in the numerical experiments in section 6, LMS is significantly faster than NUS (especially for higher-order nets) while still achieving higher-order rates of RMSE convergence.
@@
Gray code ordering of digital nets enables computing the next point ${\boldsymbol{x}}_{i+1}$ from ${\boldsymbol{x}}_{i}$ by only adding a single column from each generating matrix rather. Specifically, the $q^{\mathrm{th}}$ column of each generating matrix gets digitally added to the previous point where $q-1$ is the index of the only digit to be changed in Gray code ordering. Gray code orderings for $b=2$ and $b=3$ are shown in table 1.
@@
00 $0000_{2}$ $0000_{2}$ $00_{3}$ $00_{3}$
$1$ $0001_{2}$ $0001_{2}$ $01_{3}$ $01_{3}$
$2$ $0010_{2}$ $0011_{2}$ $02_{3}$ $02_{3}$
$3$ $0011_{2}$ $0010_{2}$ $10_{3}$ $12_{3}$
$4$ $0100_{2}$ $0110_{2}$ $11_{3}$ $11_{3}$
$5$ $0101_{2}$ $0111_{2}$ $12_{3}$ $10_{3}$
$6$ $0110_{2}$ $0101_{2}$ $20_{3}$ $20_{3}$
$7$ $0111_{2}$ $0100_{2}$ $21_{3}$ $21_{3}$
$8$ $1000_{2}$ $1000_{2}$ $22_{3}$ $22_{3}$
Table 1. Gray code order for bases $b=2$ and $b=3$. In Gray code order only one digit is incremented or decremented by $1$ (modulo $b$) when $i$ is incremented by $1$.
The following code generates a base $2$ digital net with $R$ independent LMS, $R$ independent digital shifts, and $\alpha=2$ interlacing.
@@ 00 $0000_{2}$ $0000_{2}$ $00_{3}$ $00_{3}$
randomize = "LMS_DS", # Matousek’s LMS with a digital shift
# options are [LMS_DS, NUS, DS, LMS, False]
t_lms = 64, # number of LMS bits i.e. number of rows in S_j
@@
replications = 15, # R
graycode = False, # use natural order instead of Gray code order
@@
generating_matrices = "LDData:mps.sobol_Cs.txt")
x = dnb2(2**16) # a numpy.ndarray with shape 15 x 65536 x 52
Here we have used a set of Sobol’ generating matrices from Joe and Kuo131313the “new-joe-kuo-6.21201” direction numbers from https://web.maths.unsw.edu.au/~fkuo/sobol/index.html (Joe and Kuo, 2008) which are stored in a standardized format in the LDData repository. Other generating matrices from LDData may be used by passing in a file name from https://github.com/QMCSoftware/LDData/blob/main/dnet/ or by passing an explicit array.
4.3. Halton sequences
The digital sequences described in the previous section used a fix prime base $b$. One may allow each dimension $j\in\{1,\dots,d\}$ to have its own prime base $b_{j}$. The most popular of such constructions is the Halton sequence which sets $b_{j}$ to the $j^{\mathrm{th}}$ prime, sets ${\mathsf{C}}_{j}$ to the identity matrix, and sets $t_{\mathrm{max}}=m$. This enables the simplified construction of Halton points $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ via
@@ 4.3. Halton sequences
Almost all the methods described for digital sequences are immediately applicable to Halton sequences after accounting for the differing bases across dimensions. However, digital interlacing is not generally applicable when the bases differ. Halton with random starting points has also been explored in (Wang and Hickernell, 2000), although we do not consider this here. The following code generates a Halton point set with $R$ independent LMS and $R$ independent digital permutations.
@@
randomize = "LMS_PERM", # Matousek’s LMS with a digital permutation
# options are [LMS_PERM, LMS_DS, LMS, PERM, DS, NUS, QRNG, FALSE]
t_lms = 64, # number of LMS digits i.e. number of rows in S_j
replications = 15, # R
@@
x = halton(2**10) # a numpy.ndarray with shape 15 x 1024 x 52
The QRNG randomization follows the QRNG software package (Hofert and Lemieux, 2023) in generating a generalized Halton point set (Faure and Lemieux, 2009) using a deterministic set of permutation scrambles and random digital shifts.
5. Fast transforms and kernel computations
Recall from the introduction that we would like to compute the matrix-vector product ${\mathsf{K}}{\boldsymbol{y}}$ and solve the linear system ${\mathsf{K}}^{-1}{\boldsymbol{y}}$ where ${\boldsymbol{y}}$ is some known length $n$ vector and ${\mathsf{K}}=\{K({\boldsymbol{x}}_{i},{\boldsymbol{x}}_{k})\}_{i,k=0}^{n-1}$ is a $n\times n$ symmetric positive definite (SPD) Gram matrix based on an SPD kernel $K$ and point set $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$. We may reduce the standard ${\mathcal{O}}(n^{2})$ and ${\mathcal{O}}(n^{3})$ costs of solving these respective problems to both be ${\mathcal{O}}(n\log n)$ by inducing structure into ${\mathsf{K}}$ using special point set-kernel pairings. Specifically, pairing lattice points with a shift invariance kernel yields a circulant Gram matrix ${\mathsf{K}}$ for which computations can be done quickly using the bit-reversed fast Fourier transform (FFT) and bit-reversed inverse FFT (IFFT). Similarly, pairing a digital net with digitally-shift-invariant kernel yields a nested block Toeplitz matrix for which computations can be done quickly using the fast Walsh Hadamard Transform (FWHT). Shift-invariant and digitally-shift-invariant kernels are shown in fig. 3. The reduced costs achieved by inducing these structures are summarized in table 2.
Figure 3. Shift-invariant kernels (top row) and digitally-shift-invariant kernels (bottom row). Here $\alpha$ denotes the smoothness of the kernel. The $\alpha=3$ and $\alpha=4$ digitally-shift-invariant kernels are nearly indistinguishable. Shift-invariant kernels are described in section 5.1 while digitally-shift-invariant kernels are described in section 5.2.
[…]
@@ 5. Fast transforms and kernel computations
general SPD ${\mathcal{O}}(n^{2})$ ${\mathcal{O}}(n^{3})$ ${\mathcal{O}}(n^{2})$ ${\mathcal{O}}(n^{2})$ standard
lattice SPD SI ${\mathcal{O}}(n)$ ${\mathcal{O}}(n\log n)$ ${\mathcal{O}}(n\log n)$ ${\mathcal{O}}(n\log n)$ FFT-based
digital net SPD DSI ${\mathcal{O}}(n\log n)$ ${\mathcal{O}}(n)$ ${\mathcal{O}}(n\log n)$ ${\mathcal{O}}(n\log n)$ FWHT-based
Table 2. Comparison of storage and cost requirements for kernel methods. Decomposition of the symmetric positive definite (SPD) Gram matrix ${\mathsf{K}}$ is the cost of computing the eigendecomposition or Cholesky factorization. The costs of matrix-vector multiplication and solving a linear system are the costs after performing the decomposition. Both storage and kernel computation costs are greatly reduced by pairing certain low-discrepancy point sets with special shift-invariant (SI) or digitally-shift-invariant (DSI) kernels.
The following code can be used to construct a Gram matrix with a lattice point set and shift-invariant kernel. The Gram matrix is then used perform matrix-vector multiplication and solve a linear system. Analogous code using qp.DigitalNetB2, qp.KernelDigShiftInvar, and qp.FastGramMatrixDigitalNetB2 is also supported.
d = 5 # dimension
n = 2**20 # 2^20 points
dd_obj = qp.Lattice(dimension=d, seed=7)
# qp.Lattice defaults to natural ordering with a single random shift
kernel_obj = qp.KernelShiftInvar( # matching kernel
dimension = d,
lengthscales = 0.1*np.ones(d,dtype=float), # product weights
alpha = 4*np.ones(d,dtype=int)) # smoothness parameters
K = qp.FastGramMatrixLattice( # n x n Gram matrix
kernel_obj = kernel_obj,
dd_obj = dd_obj,
n1 = n,
n2 = n)
y = np.random.rand(n) # get a random length n vector
a = K@y # compute the matrix-vector product a=Ky at O(n log n) cost
b = K.solve(y) # solve linear system Kb=y for b at O(n log n) cost
Schematics of the bit-reversed FFT and FWHT underlying these fast computations are shown in fig. 4. The one-dimensional bit-reversed FFT, one-dimensional bit-reversed IFFT, and one-dimensional FWHT are implemented into the QMCPy functions qp.fftbr, qp.ifftbr, and qp.fwht respectively. These fast transforms act along the last dimension which is required to to have length $2^{m}$ for some $m\geq 0$. The following subsections detail the matching point set-kernel pairings and provide methods to quickly update the Gram matrix product or linear system solution after doubling the point size.
@@
Figure 4. Bit-reversed FFT and FWHT schemes. The bit-reversed FFT is performed via a decimation in time algorithm without the initial reordering of inputs. The bit-reversed IFFT may be performed by propagating $\{{\widetilde{y}}_{i}\}_{i=0}^{2^{m}-1}$ to $\{y_{i}\}_{i=0}^{2^{m}-1}$ in the other direction through the bit-reversed FFT algorithm.
[…]
5.1. Shift invariant kernels, lattices, and the bit-reversed FFT/IFFT
A kernel is said to be shift-invariant when $K({\boldsymbol{u}},{\boldsymbol{v}})={\widetilde{K}}(({\boldsymbol{u}}-{\boldsymbol{v}})\mod 1)$ for some ${\widetilde{K}}$ i.e. the kernel is only a function of the component-wise difference between inputs modulo $1$. One set of shift-invariant kernels take the form
(4) $K({\boldsymbol{u}},{\boldsymbol{v}})=\sum_{{\mathfrak{u}}\subseteq{\mathcal{D}}}\gamma_{\mathfrak{u}}\prod_{j\in{\mathfrak{u}}}\eta_{\alpha_{j}}((u_{j}-v_{j})\mod 1)$

### v2

@@ 5.2 Digital nets
and ${\boldsymbol{\Delta}}_{j}$ is the $j^{\mathrm{th}}$ column of $\mathsf{\Delta}$}. Randomly shifted digital nets use ${\mathsf{\Delta}}\overset{\mathrm{IID}}{\sim}{\mathcal{U}}\{0,\dots,b-1\}^{t_{\mathrm{max}}\times d}$, i.e., each digit is chosen uniformly from $\{0,\dots,b-1\}$.
Digitally permuted digital net
In what follows we will denote permutations of $\{0,\dots,b-1\}$ by $\pi$. Suppose we are given a set of permutations
@@
Then we may construct the digitally permuted digital net $\{{\boldsymbol{x}}_{i}\}_{i=0}^{b^{m}-1}$ where
@@
Nested Uniform Scrambling (NUS)
NUS is often called Owen scrambling for its conception in [Owen, 1995]. As before, $\pi$ denotes permutations of $\{0,\dots,b-1\}$. Now suppose we are given a set of permutations
@@
Then a nested uniform scrambling of a digital net is $\{{\boldsymbol{x}}_{i}\}_{i=0}^{b^{m}-1}$ where
@@
our implementation cannot afford to generate all permutations a priori. Instead, permutations are generated and stored only as needed. As with digitally-permuted digital nets, NUS uses independent uniform random permutations.
The randomization routines described above are visualized in Figure 2.
@@
for $j=1,\dots,d$. Following [Owen, 2003], let us denote elements in $\{1,\dots,b-1\}$ by $h$ and elements in $\{0,\dots,b-1\}$ by $g$. Then common structures for ${\mathsf{S}}_{j}$ include
@@
which corresponds to Matoušek’s linear scrambling [Matoušek, 1998], Tezuka’s $i$-binomial scrambling [Tezuka, 2002], and Owen’s striped LMS [Owen, 2003] (not to be confused with NUS which is often called Owen scrambling). Random LMS chooses $g$ and $h$ values all independently and uniformly from $\{1,\dots,b-1\}$ and $\{0,\dots,b-1\}$ respectively. Owen [2003] analyzes these scramblings and their connection to NUS.
Digital interlacing enables the construction of higher-order digital nets. For integer order $\alpha\geq 1$, interlacing ${\mathsf{A}}_{1},{\mathsf{A}}_{2},\dots\in\{0,\dots,b\}^{t_{\mathrm{max}}\times m}$ gives $\widehat{{\mathsf{A}}}_{1},\widehat{{\mathsf{A}}}_{2},\dots\in\{0,\dots,b\}^{\alpha t_{\mathrm{max}}\times m}$ satisfying $\widehat{{\mathsf{A}}}_{jtk}=A_{\widehat{j},\widehat{t},k}$ where $\widehat{j}=\alpha(j-1)+(t\mod\alpha)+1$ and $\widehat{t}=\lfloor t/\alpha\rfloor$ for $j\geq 1$ and $0\leq t<\alpha t_{\mathrm{max}}$ and $1\leq k\leq m$. For example, with $m=2$, $t_{\mathrm{max}}=2$, and $\alpha=2$ we have
${\mathsf{A}}_{1}=\begin{pmatrix}a_{101}&a_{102}\\ a_{111}&a_{112}\end{pmatrix}\qquad{\mathsf{A}}_{2}=\begin{pmatrix}a_{201}&a_{202}\\ a_{211}&a_{212}\end{pmatrix}\qquad{\mathsf{A}}_{3}=\begin{pmatrix}a_{301}&a_{302}\\ a_{311}&a_{312}\end{pmatrix}\qquad{\mathsf{A}}_{4}=\begin{pmatrix}a_{401}&a_{402}\\ a_{411}&a_{412}\end{pmatrix}\qquad\dots$
$\widehat{{\mathsf{A}}}_{1}=\begin{pmatrix}a_{101}&a_{102}\\ a_{201}&a_{202}\\ a_{111}&a_{112}\\ a_{211}&a_{212}\end{pmatrix}\qquad\widehat{{\mathsf{A}}}_{2}=\begin{pmatrix}a_{301}&a_{302}\\ a_{401}&a_{402}\\ a_{311}&a_{312}\\ a_{411}&a_{412}\end{pmatrix}\qquad\dots$
Without scrambling, higher-order digital nets may be directly generated from $\widehat{{\mathsf{C}}}_{1},\dots,\widehat{{\mathsf{C}}}_{d}$, the interlaced generating matrices resulting from interlacing ${\mathsf{C}}_{1},\dots,{\mathsf{C}}_{\alpha d}$. Higher-order NUS [Dick, 2011] requires generating a digital net from ${\mathsf{C}}_{1},\dots,{\mathsf{C}}_{\alpha d}$, applying NUS to the resulting $\alpha d$ dimensional digital net, and then interlacing the resulting digit matrices $\{{\boldsymbol{\mathsf{z}}}_{i,1}^{\intercal}\}_{i=0}^{b^{m}-1},\dots,\{{\boldsymbol{\mathsf{z}}}_{i,\alpha d}^{\intercal}\}_{i=0}^{b^{m}-1}\in\{0,\dots,b\}^{t_{\mathrm{max}}\times m}$. For higher-order LMS, one applies LMS to the generating matrices ${\mathsf{C}}_{1},\dots,{\mathsf{C}}_{\alpha d}$ to get ${\widetilde{\mathsf{C}}}_{1},\dots,{\widetilde{\mathsf{C}}}_{\alpha d}$, then interlaces ${\widetilde{\mathsf{C}}}_{1},\dots,{\widetilde{\mathsf{C}}}_{\alpha d}$ to get $\widehat{{\widetilde{\mathsf{C}}}}_{1},\dots,\widehat{{\widetilde{\mathsf{C}}}}_{d}$, then generates the digital net from $\widehat{{\widetilde{\mathsf{C}}}}_{1},\dots,\widehat{{\widetilde{\mathsf{C}}}}_{d}$. As we show in the numerical experiments in Section 7, LMS is significantly faster than NUS (especially for higher-order nets) while still achieving higher-order rates of RMSE convergence.
@@
Gray code ordering of digital nets enables computing the next point ${\boldsymbol{x}}_{i+1}$ from ${\boldsymbol{x}}_{i}$ by only adding a single column from each generating matrix. Specifically, the $q^{\mathrm{th}}$ column of each generating matrix gets digitally added to the previous point where q1q-1 is the index of the only digit to be changed in Gray code ordering. Gray code orderings for $b=2$ and $b=3$ are shown in Table 1.
@@
0 $00_{2} 0000$$0_{2} 00$$_{3} 00$$0_{3}$
11 $001_{2} 00$$001_{2}$ $01_{3}$ $01_{3}$
22 $010_{2} 00$$011_{2}$ $02_{3}$ $02_{3}$
33 $011_{2} 00$$010_{2}$ $10_{3}$ $12_{3}$
44 $00_{2} 0110$$0_{2} 11$$_{3} 11$$1_{3}$
55 $0101_{2}$ $0111_{2}$ $12_{3}$ $10_{3}$
66 $10_{2} 0101$$01_{2}$ $20_{3}$ $20_{3}$
77 $0111_{2}$ $0100_{2}$ $21_{3}$ $21_{3}$
88 $1000_{2}$ $1000_{2}$ $2_{3} 22$$2_{3}$
Table 1: Gray code order for bases $b=2$ and $b=3$. In Gray code order only one digit is incremented or decremented by 11 (modulo $b$) when $i$ is incremented by $1$.
The following code generates $\alpha=2$ higher-order digital nets in base $2$ with $R$ independent LMS and digital shift combinations.
Listing 3: Generate randomized higher-order digital nets.
@@ 88 $1000_{2}$ $1000_{2}$ $2_{3} 22$$2_{3}$
randomize = "LMS DS", # Matousek’s LMS then a digital shift
# other options ["NUS", "DS", "LMS", None]
t = 64, # number of LMS bits, i.e., number of rows in S_j
@@
replications = 16, # R
order = "radical inverse", # also supports "Gray code"
@@
generating_matrices = "joe_kuo.6.21201.txt")
x = dnb2(2**16) # a numpy.ndarray with shape (16, 65536, 52)
Here we have used a set of Sobol’ generating matrices from Joe and Kuo161616“new-joe-kuo-6.21201” direction numbers from https://web.maths.unsw.edu.au/~fkuo/sobol/index.html [Joe and Kuo, 2008] which are stored in a standardized format in the LDData repository. Other generating matrices from LDData may be used by passing in a file name from https://github.com/QMCSoftware/LDData/blob/main/dnet/ or by passing an explicit array.
5.3 Halton sequences
The digital sequences described in the previous section used a fixed prime base $b$. One may allow each dimension $j\in\{1,\dots,d\}$ to have its own prime base $b_{j}$. The most popular of such constructions is the Halton sequence which sets $b_{j}$ to the $j^{\mathrm{th}}$ prime, sets ${\mathsf{C}}_{j}$ to the identity matrix, and sets $t_{\mathrm{max}}=m$. This enables the simplified construction of Halton points $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$ via
@@ 5.3 Halton sequences
Almost all the methods described for digital sequences are immediately applicable to Halton sequences after accounting for the differing bases across dimensions. However, digital interlacing is not generally applicable when the bases differ. Halton with random starting points has also been explored in [Wang and Hickernell, 2000], although we do not consider this here. The following code generates a Halton point set with $R$ independent LMS and digital permutation combinations. Specifically, we generate $R$ independent LMS-Halton point sets and then apply an independent permutation scramble to each.
Listing 4: Generate randomized Halton points.
@@
randomize = "LMS PERM", # Matousek’s LMS then a digital perm
# or ["LMS DS", "LMS", "PERM", "DS", "NUS", "QRNG", None]
t = 64, # number of LMS digits, i.e., number of rows in S_j
replications = 16, # R
@@
x = halton(2**10) # a numpy.ndarray with shape (16, 1024, 52)
The "QRNG" randomization follows the QRNG software package [Hofert and Lemieux, 2023] in generating a generalized Halton point set [Faure and Lemieux, 2009] using a deterministic set of permutation scrambles and random digital shifts.
6 Fast transforms and kernel computations
Recall from the motivation in Section 3.2 that we would like to compute the matrix-vector product ${\mathsf{K}}{\boldsymbol{y}}$ and solve the linear system ${\mathsf{K}}^{-1}{\boldsymbol{y}}$ where $\boldsymbol{y}$} is some known length $n$ vector and ${\mathsf{K}}=\{K({\boldsymbol{x}}_{i},{\boldsymbol{x}}_{i^{\prime}})\}_{i,i^{\prime}=0}^{n-1}$ is an $n\times n$ symmetric positive definite (SPD) Gram matrix based on a SPD kernel $K$ and point set $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}$. Table 2 compares our fast kernel methods against standard techniques in terms of both storage and computational costs. Figure 3 gives schematics of the FFTBR and FWHT algorithms which enable these fast kernel methods. The following subsections detail these special kernel-point set pairings and how the structure induce in $\mathsf{K}$} can be exploited for fast kernel computations.
@@ 6 Fast transforms and kernel computations
any general SPD ${\mathcal{O}}(n^{2})$ ${\mathcal{O}}(n^{3})$ ${\mathcal{O}}(n^{2})$ ${\mathcal{O}}(n^{2})$ Cholesky
rank-1 lattice SPD SI ${\mathcal{O}}(n)$ ${\mathcal{O}}(n\log n)$ ${\mathcal{O}}(n\log n)$ ${\mathcal{O}}(n\log n)$ FFTBR/IFFTBR
digital net SPD DSI ${\mathcal{O}}(n)$ ${\mathcal{O}}(n\log n)$ ${\mathcal{O}}(n\log n)$ ${\mathcal{O}}(n\log n)$ FWHT
Table 2: Comparison of storage and cost requirements for kernel methods. Decomposition of the symmetric positive definite (SPD) Gram matrix $\mathsf{K}$} is the cost of computing the eigendecomposition or Cholesky factorization. The costs of matrix-vector multiplication and solving a linear system are the costs after performing the decomposition. Both storage and kernel computation costs are greatly reduced by pairing certain low-discrepancy point sets with special shift-invariant (SI) or digitally-shift-invariant (DSI) kernels. These fast algorithms rely on the fast Fourier transform in bit-reversed order (FFTBR), its inverse (IFFTBR), and the fast Walsh–Hadamard transform (FWHT).
@@
Figure 3: Bit-reversed FFT and FWHT schemes. The bit-reversed FFT is performed via a decimation in time algorithm without the initial reordering of inputs. The bit-reversed IFFT may be performed by propagating $\{{\widetilde{y}}_{i}\}_{i=0}^{2^{m}-1}$ to $\{y_{i}\}_{i=0}^{2^{m}-1}$ in the other direction through the bit-reversed FFT algorithm.
6.1 SI kernels, rank-1 lattices, and the FFTBR/IFFTBR
A kernel is said to be shift-invariant (SI) when
$K({\boldsymbol{x}},{\boldsymbol{x}}^{\prime})={\widetilde{K}}(({\boldsymbol{x}}-{\boldsymbol{x}}^{\prime})\mod 1)$
for some $\widetilde{K}$}, i.e., the kernel is only a function of the component-wise difference between inputs modulo $1$. One set of SI kernels take the form
$K({\boldsymbol{x}},{\boldsymbol{x}}^{\prime})=\gamma\sum_{{\mathfrak{u}}\subseteq\{1,\dots,d\}}\eta_{\mathfrak{u}}\prod_{j\in{\mathfrak{u}}}K_{\alpha_{j}}(x_{j},x_{j}^{\prime})$ (5)